# Fréchet Power-Scenario Distance: A Metric for Evaluating Generative AI Models across Multiple Time-Scales in Smart Grids

### v1

@@ II-C Fréchet Power-Scenario Distance (FPD) calculation
Once robust representations are obtained from the feature extractor, we can then compare the distribution similarity between the two datasets in the feature space. Three distribution evaluation metrics have often been adopted[7]: (1) KL-divergence, (2) Kernel Distance (Maximum Mean Discrepancy - MMD), and (3) Fréchet Distance (Wasserstein-2 Distance).
The features can often be approximately viewed as having a Gaussian distribution. Denote the feature vectors extracted from two sets $x_{r}^{1}$ and $x_{r}^{2}$ at resolution $r$ as $z_{r}^{1}$ and $z_{r}^{2}$. The means of extracted features can then be expressed as:
@@
Using the KL-divergence as the evaluating metric, the similarity between $z_{r}^{1}$ and $z_{r}^{2}$ can be assessed as [15]:
$\text{KL}(z_{r}^{1}||z_{r}^{2})=$ $\frac{1}{2}\bigg{(}\text{Tr}(\Sigma_{2}^{-1}\Sigma_{1})+(\mathbf{m}_{2}-\mathbf{m}_{1})^{T}\Sigma_{2}^{-1}(\mathbf{m}_{2}-\mathbf{m}_{1})$ (5)
$-d+\log\left(\frac{\det(\Sigma_{2})}{\det(\Sigma_{1})}\right)\bigg{)},$
where $d$ is the dimensionality of the feature space $z_{r}^{1}$ and $z_{r}^{2}$. It is important to note that when applying the KL-divergence to compare the distance between two normal distributions, both $z_{r}^{1}$ and $z_{r}^{2}$ must have the same dimensionality. Because our goal is to design a metric that can compare datasets in different resolutions which have different feature dimensions, the KL-divergence is not suitable for our purpose. It is also well known that the KL divergence has the disadvantage of being asymmetric and unstable in some cases.
@@
where $k(x,y)$ is a kernel function (e.g., Gaussian kernel). Choosing a proper kernel is extremely important for the metric to be effective, as different kernel functions can lead to varying metric performance.
We propose to use the Fréchet Distance due to its stability and consistency as a metric based on our experiments. Since this metric is specifically tailored to evaluate power system scenarios, we name it Fréchet Power-Scenario Distance (FPD). The $\text{FPD}(z_{r}^{1},z_{r}^{2})$ between the extracted features at resolution $r$ is calculated as:
@@
FPD is essentially a form of the Fréchet Distance, where we assume that the features of the two datasets follow Gaussian distributions. This metric quantifies the similarity between the two distributions: a smaller FPD value indicates a more significant similarity between the feature distributions, ultimately reflecting a higher similarity between the datasets.
@@ III Model Training
The training data used in this study consists of multiple open-source datasets, each capturing different generative tasks. These datasets include various types of demand, EV charging profiles, and various types of electricity generation. Each dataset varies in terms of resolution and timescale, contributing to a comprehensive representation of the power system domain. The specific characteristics of each dataset are summarized in Table II.
TABLE II: Summary of Datasets Used for Training
Dataset Description Source Resolution Timescale
1 Wind Generation[16] NREL 5-minute 1 year
2 Solar Generation[17] NREL 5-minute 1 year
3 EV charging profiles[18] Northwest Power and Conservation Council Hourly 1 year
4 Commercial user load[19] OpenEI.org Hourly 1 years
5 Residential user load[19] OpenEI.org Hourly 1 years
6 Load Data in Texas[20] ERCOT Hourly 3 years
7 Demand by balancing authority[21] EIA daily 3 years
8 Generation by balancing authority[21] EIA daily 3 years
9 Utility Scale Electricity Net Generation[22] EIA monthly 14 years
10 Phasor Measurement Unit PSSE simulations 0.016-second(60Hz) 50 hours
For steady-state model, each module in the feature extraction model is trained individually and sequentially from the bottom level (5-min module) to the top level (yearly module). Once a certain level module has finished training, the parameters will be fixed, and we start to train the next level module. The training data for the module at each resolution is $I_{r_{s}}$ The training dataset is shown in Table II
@@ 9 Utility Scale Electricity Net Generation[22] EIA monthly 14 years
The first term in the (8), $\sum_{k=1}^{K}\left(f_{\text{reg}}^{k}(I_{r_{s}})-y_{i}^{\text{reg}}\right)^{2}$, corresponds to the Mean Squared Error (MSE) loss across a total of $K=9$ regression tasks. Each regression task is designed for leading extractors to capture features that reflect a specific statistical characteristic of the data, represented by $y_{i}^{\text{reg}}$. The list of regression tasks is shown in Table V in Appendix. The second term,$-\sum\limits_{i=1}^{C}y_{i}^{\text{clf}}\log\left(\text{softmax}(f_{\text{clf}}(I_{r_{s}}))\right)$, represents the Cross-Entropy Loss for the classification task which categorize data based on classification labels $y_{i}^{\text{clf}}$. $y_{i}^{\text{clf}}$ is determined by the dataset index shown in Table II and $C$ is the number of categories of dataset. The hyper-parameters for training are shown in Table VI in Appendix.
For transient-state model training, we employ the fault type, amplitude minimum, and maximum as supervised pretraining labels. We then exploit the pre-trained feature extractor to evaluate synthetic samples. The length of a data sequence is 960, and input data is encoded into a 2048-dimension representation. Moreover, the hyper-parameters for training are shown in Table VII in Appendix. Once training is complete, the tasks are irrelevant, and the obtained module $\mathcal{M}(I_{r_{s}})$ will be used for feature extraction.
@@ IV Effectiveness Evaluation of the FPD
IV-A Effectiveness on different disturbances
(a) Gaussian Noise
(b) Missing Data
(c) Solar Data Contaminate
(d) Gaussian Smooth
(e) Error Cumulate
(f) Time Shift
@@ IV-A Effectiveness on different disturbances
Figure 2: Effect of six different types of disturbances on the dataset $X$: (2(a)) Gaussian Noise is added with mean 0 and varying variances $\alpha=[0,0.16,1.6,4]$, where a larger $\alpha$ indicates more noise. (2(b)) Missing Data is simulated by randomly replacing $\alpha\times 100\%$ of the points in each sample of $X$ with zeros, where $\alpha=[0,0.1,0.25,0.5]$. (2(c)) Solar Data Contamination replaces $\alpha\times 100\%$ of the points in $X$ with corresponding points from a solar power dataset $Y$[17], where $\alpha=[0,0.25,0.5,0.75]$. (2(d)) Gaussian Smooth applies Gaussian filter which has variance $\alpha=[0,10,20,30]$ to the original dataset. Where a larger $\alpha$ indicates more significant smoothing. (2(e)) Error Cumulate is simulated by adding an error, which is times with a random Gaussian noise with a mean of 1 and varying variances $\alpha=[0,0.005,0.01,0.03]$ at each time-step. (2(f)) Time Shift is a potential problem that exists in time-series generation. To simulate this effect, we shift the data forward by $\alpha=[0,40,60,80]$ intervals.
@@
In this section, we conduct several experiments to demonstrate the effectiveness of FPD. The experiments are designed based on the Wind Generation dataset 1 in Table II. We progressively increase the disturbance applied to the original dataset X in each experiment to provide a controllable and scalable measure of FPD performance.
@@
  • Gaussian Noise: In this case, Gaussian noise with a mean of 0 and a variance $\alpha$ is added to $X$. Where $\alpha\in[0,0.16,1.6,4]$ represents the disturbance added. Larger $\alpha$ indicates more noise added.
@@
  • Solar Data Contamination: To show FPD can distinguish the data from another dataset, we contaminate the wind dataset $X$ with data from the solar power dataset $Y$ [17]. The disturbance $\alpha\in[0,0.25,0.5,0.75]$ indicates that $\alpha\times 100\%$ of the samples in $X$ are replaced by corresponding points from $Y$.
@@
  • Time Shift: Time shift in time-series data is a common issue in generative models, often leading to misalignment of temporal features between the original and generated data. This misalignment can be particularly problematic in power system data, where certain temporal features, such as peaks, carry critical information. To simulate this, we shift the data forward by intervals of $\alpha=[0,40,60,80]$, creating misleading temporal dependencies. Our results demonstrate that FPD effectively detects and captures these temporal shifts.
As shown in Fig. 2, KL-divergence can be inconsistent due to its asymmetry feature, while Kernel Distance may fail to capture significant changes and is sensitive to kernel choice. Additionally, both methods also exhibit scale variability across disturbances. In contrast, Fréchet Distance consistently reflects disturbance levels and maintains stable scaling, making it a more reliable evaluation metric. For these reasons, we select Fréchet Distance as the evaluation metric and refer to our method as Fréchet Power-Scenario Distance (FPD).
@@
IV-B Comparison of FPD with conventional evaluation metrics
Figure 4: Data-level Fréchet distance (FD) VS. FPD
IV-B1 Sample-Wise Euclidean Distance (MAPE) vs. Distribution-Wise Distance (FPD)
@@ IV-B Comparison of FPD with conventional evaluation metrics
Traditional Euclidean-based metrics such as Mean Absolute Percentage Error (MAPE) typically perform sample-wise comparisons, making the results highly dependent on the selection of sample pairs between the two datasets. This approach can lead to inconsistent and potentially invalid outcomes. To illustrate this, we revisit the Gaussian noise case presented in Fig. 2(a). In this example, we make MAPE to match sample pairs randomly between the original and noise-perturbed datasets. As shown in Fig. 3, in some instances, higher noise levels result in lower MAPE values, which contradicts the ground truth. In contrast, the proposed FPD metric demonstrates expected behavior, providing a more effective measurement.
IV-B2 Data-Level Distribution Comparison (Fréchet Distance) vs. Feature-Level Distribution Comparison (FPD)
As previously noted, the data-level Fréchet Distance (FD) metric is often ineffective at capturing high-level complex patterns. FPD metric addresses this issue by incorporating feature extraction and measuring the similarity between distributions in the feature space. Feature extraction enables a better representation of higher-level patterns and structures within the data, allowing FPD to capture more relevant similarities for evaluating the quality of generated data.
To demonstrate this advantage, we constructed a fabricated dataset by randomly sampling from a distribution with a similar mean and covariance to the original data. Despite these similarities, the fabricated dataset lacks the temporal patterns present in the original dataset, resulting in a sparse relation. As shown in Fig. 4, while FD fails to identify this disparity, FPD accurately reflects the weaker relationship between the fabricated and original datasets.
IV-C Coherence between FPD and the downstream tasks
To assess whether FPD’s indication of high-quality data can be used to determine the practical usefulness for downstream tasks, we further evaluate its efficacy for transient-state case using PMU data containing three types of faults. We employ the Brownian noises to mimic the accumulating temporary errors in the PMU data prediction as depicted in Fig. 5. As shown in Table III, we cross-evaluate different fault scenarios. Hence, we can observe non-zero diagonal elements in the FPD score matrix, along with the diagonal values being significantly lower than the non-diagonal elements, showing the capability of identifying different scenarios other than evaluating generation quality. To further prove the capability of FPD can handle the downstream task in steady-state case, we reuse the wind dataset[16] and define the sub-task as data classification based on the average power across the day as used in Chen’s work [23]. The three cases differ based on average power levels, as shown in Fig. 6. Moreover, the cross-evaluation of different mean value scenarios is shown in the Table.IV again highlights the alignment between the FPD evaluation and the specific sub-task.
Figure 5: Ground-truth and synthetic PMU voltage magnitude curves under different fault scenarios. Synthetic data is modulated using Brownian noises.
TABLE III: Cross-evaluation on FPD performance for PMU data.
Synthetic $|V|$ Ground-truth $|V|$ bus_trip branch_trip branch_fault
bus_trip 744.26 1086.07 2586.15
branch_trip 1462.24 846.03 2720.04
branch_fault 2516.35 2401.13 687.99
Figure 6: Ground-truth and synthetic wind power under different mean scenarios. Synthetic data is modulated using Brownian noises.
TABLE IV: Cross-evaluation on FPD performance for wind power data.
Synthetic $|MW|$ Ground-truth $|MW|$ $\bar{x}<0.5$ $\bar{x}<3$ $\bar{x}>6$
$\bar{x}<0.5$ 736.97 1300.42 901.36
$\bar{x}<3$ 1300.42 565.35 2720.04
$\bar{x}>6$ 901.36 989.83 686.51
IV-D Cross-model comparison with FPD

### v2

@@ II-C Fréchet Power-Scenario Distance (FPD) calculation
Once robust representations are obtained from the feature extractor, we can then compare the distribution similarity between the two datasets in the feature space. Three distribution-level evaluation metrics are commonly used[20]: (1) Jensen–Shannon (JS) Divergence, (2) Kernel Distance (Maximum Mean Discrepancy - MMD), and (3) Fréchet Distance (Wasserstein-2 Distance).
Denote the feature vectors extracted from two sets $x_{r}^{1}$ and $x_{r}^{2}$ at resolution $r$ as $z_{r}^{1}$ and $z_{r}^{2}$. The means of extracted features can then be expressed as:
@@
Using the Jensen–Shannon (JS) divergence—the symmetrized and smoothed variant of KL [21] —the similarity between $z_{r}^{1}$ and $z_{r}^{2}$ is defined as
$\mathrm{JS}(z_{r}^{1}\,\|\,z_{r}^{2})$ $=\tfrac{1}{2}\,\mathrm{KL}\!\big(\mathcal{N}(\mathbf{m}_{1},\Sigma_{1})\,\big\|\,\mathcal{N}(\mathcal{M}\big)$ (5)
$\quad+\tfrac{1}{2}\,\mathrm{KL}\!\big(\mathcal{N}(\mathbf{m}_{2},\Sigma_{2})\,\big\|\,\mathcal{N}(\mathcal{M}\big).$
where $\mathcal{M}$ is the (Gaussian) mixture approximation with mean and covariance $\mathbf{m}_{M}=\tfrac{1}{2}(\mathbf{m}_{1}+\mathbf{m}_{2})$ and $\Sigma_{M}=\tfrac{1}{2}(\Sigma_{1}+\Sigma_{2})$, and $\mathrm{KL}(\cdot\|\cdot)$ denotes the closed-form KL divergence between Gaussians:
$\mathrm{KL}\!\left(\mathcal{N}(\mathbf{m}_{1},\Sigma_{1})\,\middle\|\,\mathcal{N}(\mathbf{m}_{2},\Sigma_{2})\right)=\tfrac{1}{2}\Big[\operatorname{tr}(\Sigma_{2}^{-1}\Sigma_{1})+$ (6)
$(\mathbf{m}_{2}-\mathbf{m}_{1})^{\top}\Sigma_{2}^{-1}(\mathbf{m}_{2}-\mathbf{m}_{1})-d+\log\!\frac{\det\Sigma_{2}}{\det\Sigma_{1}}\Big].$
Here $d$ is the dimensionality of the feature space for $z_{r}^{1}$ and $z_{r}^{2}$. JS divergence is symmetric and bounded, offering improved stability over plain KL. However, when the distributions have little or no support overlap, JS becomes insensitive—it can saturate at its maximum and fail to track changes in similarity reliably, as illustrated in [22].
@@
where $k(x,y)$ is a kernel function (e.g., Gaussian kernel). Choosing a proper kernel is extremely important for the metric to be effective, as different kernel functions can lead to varying metric performances.
We use the Fréchet distance due to its empirically observed stability and consistency. Since this metric is specifically tailored to evaluate power system scenarios, we refer to this as Fréchet Power-Scenario Distance (FPD). The $\text{FPD}(z_{r}^{1},z_{r}^{2})$ between the extracted features at resolution $r$ is calculated as:
@@
FPD is essentially a form of the Fréchet distance, where we assume that the features of the two datasets follow Gaussian distributions [^1] For practical reasons, we consider only the mean and variance of distributions. Since the Gaussian maximizes entropy subject to fixed mean and covariance, we model feature distributions as multivariate Gaussian.. This metric quantifies the similarity between the two distributions: a smaller FPD value indicates a more significant similarity between the feature distributions, ultimately reflecting a higher similarity between the datasets.
@@ III Model Training
The purpose of training is to guide the feature extraction model to produce features that are tightly coupled to the smart grid datasets while retaining strong representational power. The training data used in this study consists of multiple open-source datasets, each capturing different generative tasks. These datasets include various types of demand, EV charging profiles, and various types of electricity generation. Each dataset varies in terms of resolution and timescale, contributing to a comprehensive representation of the smart grid domain. The specific characteristics of each dataset are summarized in SI Sec.S3; Table S8.
For the steady-state model, each module in the feature extraction model is trained individually and sequentially from the bottom level (5-minute module) to the top level (yearly module). Once a certain level module has finished training, the parameters will be fixed, and we will start to train the next level module. The training data for the module at each resolution is $I_{r_{s}}$. The training dataset is shown in SI Sec.S3; Table S8.
@@
[…]
For transient-state model training, we use fault type, minimum amplitude, and maximum amplitude as supervised pretraining labels. We then exploit the pretrained feature extractor to evaluate synthetic samples. The length of a data sequence is 960, and the input data is encoded into a 2048-dimensional representation. Moreover, the hyper-parameters for training are shown in SI Sec.S1; Table S3. Once training is complete, we discard the task heads and use the trained module $\mathcal{M}(I_{r_{s}})$ solely for feature extraction.
@@ IV Effectiveness Evaluation of the FPD
IV-A Power-Aware Feature Extraction vs. Generic Encoders
In this section, we demonstrate that the proposed extractor produces representations that align more closely with these physical principles than generic time-series encoders (e.g., TS2Vec [18], TimeMixer [19]) on solar generation data. For each extractor/encoder, we compute similarity using the same Fréchet distance in the respective feature space:
  • Period Offset: Power signals exhibit strong periodicity (e.g., daily/seasonal load patterns; solar generation following the diurnal cycle—zero at night, peak near noon). To test sensitivity to disturbed periodic structure, we apply a time offset of the solar trace by $\alpha\in\{0,2,4\}$ hours. Here, $\alpha$ quantifies how far the shifted solar profile is displaced from its true period. A robust extractor/encoder should distinguish these shifted variants from the nominal series.
  • Nighttime Generation (Physical Violation): Power signals must respect physical constraints (e.g., bounded ramp rates; no solar output at night). We inject nonzero solar generation during nighttime to create a controlled violation and evaluate whether extractors/encoders can detect it. The violation level is parameterized by $\alpha\in\{0,2,3\}$, where $\alpha$ denotes the number of nighttime hours forced to have nonzero solar output.
  • Cross–Time-Resolution Consistency: A key property of power signals is that the same underlying process may be represented at different sampling resolutions; representations derived from hourly and 5-minute data should remain consistent. Our feature extractor employs a hierarchical, multi-scale design to capture structure across resolutions, whereas generic encoders (e.g., single-scale inputs) may treat different resolutions as unrelated sequences. To demonstrate this advantage, we compare features from hourly solar/wind datasets against their 5-minute counterparts and assess their similarity.
As shown in Fig. 2, the proposed extractor consistently and effectively distinguishes period shifts, detects physical feasibility violations, and preserves similarity judgments across time resolutions. By contrast, features from generic encoders do not reliably reveal phase misalignment or feasibility violations and yield inconsistent cross-resolution similarity, indicating reduced sensitivity to power-specific structure.
These findings demonstrate that our extractor is better aligned with power-system characteristics than generic models. By faithfully capturing the distinct temporal regularities, physical constraints, and multi-resolution structure of power data, it provides a solid foundation for more reliable evaluation and analysis in subsequent sections.
Figure 2: Features from the proposed power-aware extractor vs. generic time-series encoders, evaluated using the Fréchet distance in feature space in case: (a) Period Offset: solar series are shifted by $\alpha\in\{0,2,4\}$ hours. (b) Solar Physical Violation: Unrealistic power generation is injected for $\alpha\in\{0,2,3\}$ hours of solar generation during night. (c) Cross Time Resolution: 5-minute solar compared against hourly solar/wind.
IV-B Distributional Metric Suitability in Learned Feature Space
@@
Figure 3: Effectiveness of FPD with six different types of disturbances on the dataset $X$: (a) Gaussian Noise is added with mean 0 and varying variances $\alpha=[0,0.16,1.6,4]$, where a larger $\alpha$ indicates more noise. (b) Missing Data is simulated by randomly replacing $\alpha\times 100\%$ of the points in each sample of $X$ with zeros, where $\alpha=[0,0.1,0.25,0.5]$. (c) Solar Data Contamination replaces $\alpha\times 100\%$ of the points in $X$ with corresponding points from a solar power dataset YY[23], where $\alpha=[0,0.25,0.5,0.75]$. (d) Gaussian Smooth applies Gaussian filter which has variance $\alpha=[0,10,20,30]$ to the original dataset. A larger $\alpha$ indicates more significant smoothing. (e) Error Cumulate is simulated by adding an error, which is multipled by a random Gaussian noise with a mean of 1 and varying variances $\alpha=[0,0.005,0.01,0.03]$ at each time step. (f) Time Shift is a potential problem that exists in time-series generation. To simulate this effect, we shift the data forward by $\alpha=[0,40,60,80]$ intervals.
@@
In the previous section, we showed that the learned feature space produced by our extractor aligns with key smart grid properties. Here, we further present experiments that motivate the choice of the Fréchet distance over alternative distributional metrics for measuring similarity within this learned space. The experiments are designed based on the Wind Generation Dataset (Dataset 1 in SI Sec.S3; Table S8) We progressively increase the disturbance applied to the original dataset X in each experiment to provide a controllable and scalable measure of FPD performance.
@@
  • Gaussian Noise: In this case, Gaussian noise with a mean of 0 and a variance $\alpha$ is added to $X$, where $\alpha\in[0,0.16,1.6,4]$ denotes the disturbance level. A larger $\alpha$ indicates more noise added.
@@
  • Solar Data Contamination: To test cross-dataset discrimination, we contaminate the wind dataset $X$ with samples from the solar dataset $Y$ [23]. The disturbance $\alpha\in[0,0.25,0.5,0.75]$ indicates that $\alpha\times 100\%$ of the samples in $X$ are replaced by corresponding points from $Y$.
@@
  • Time Shift: Time shift in time-series data is a common issue in generative models, often leading to misalignment of temporal features between the original and generated data. This misalignment can be particularly problematic in smart grid data, where certain temporal features, such as peaks, carry critical information. To simulate this, we shift the data forward by intervals of $\alpha=[0,40,60,80]$, creating misleading temporal dependencies. Our results demonstrate that FPD effectively detects and captures these temporal shifts.
As shown in Fig. 3, JS-divergence is neither consistent nor effective across different disturbance events, while Kernel distance may fail to capture significant changes and is sensitive to kernel choice. Additionally, both methods also exhibit scale variability across disturbances. In contrast, Fréchet distance consistently reflects disturbance levels and maintains stable scaling, making it a more reliable evaluation metric. For these reasons, we select Fréchet distance as the evaluation metric and refer to our method as Fréchet Power-Scenario Distance (FPD).
IV-C Comparison of FPD with conventional evaluation metrics
@@
Figure 5: Data-level distribution metrics vs. FPD
IV-C1 Sample-Wise Euclidean Distance (MAPE) vs. Distribution-Wise Distance (FPD)
@@
Traditional Euclidean-based metrics, such as Mean Absolute Percentage Error (MAPE), typically perform sample-wise comparisons, making the results highly dependent on the selection of sample pairs between the two datasets. This approach can lead to inconsistent and potentially invalid outcomes. To illustrate this, we revisit the Gaussian noise case presented in Fig. 3. In this example, we compute MAPE after randomly matching sample pairs between the original and noise-perturbed datasets. As shown in Fig. 4, in some instances, higher noise levels result in lower MAPE values, which contradicts the ground truth. In contrast, the proposed FPD metric demonstrates expected behavior, providing a more effective measurement.
IV-C2 Data-Level Distribution Comparison vs. Feature-Level Distribution Comparison (FPD)
Although distribution-level scoring rules such as CRPS and the Energy Score, as well as distributional metrics like JS-divergence, MMD, and Wasserstein distance, avoid the sample-wise limitations discussed above, they still struggle to detect changes in generation quality when applied directly to raw data. In particular, such data-level metrics often underemphasize higher-order structure—such as temporal dependencies and cross-feature interactions.
The proposed FPD metric addresses this gap by introducing a feature-extraction stage prior to distributional comparison: data are first mapped into a learned representation, and similarity is then evaluated in this feature space. This feature-based approach highlights higher-level patterns and structures, yielding similarity estimates that more faithfully capture the quality of generated data.
To illustrate, we construct a fabricated dataset by randomly sampling from a distribution with mean and covariance matched to the original data. Despite this alignment in first- and second-order statistics, the fabricated dataset lacks the temporal structure of the original, resulting in weaker dependencies. As shown in Fig. 5, CRPS and the Energy Score show little separation between the fabricated dataset and a lightly perturbed Gaussian baseline, while the Fréchet distance also fails to flag the discrepancy. By contrast, FPD correctly identifies a weaker relationship between the fabricated and original datasets.
IV-D Robustness of FPD to Labeling Schemes and Sample Budgets for Downstream Tasks
We ask whether FPD’s indication of data quality reliably predicts downstream utility under realistic variation in labeling and sampling. In practice, given a downstream task, generative models can synthesize effectively unbounded scenario sets whose usefulness depends on both the scenario definition and the sampling strategy. We therefore test whether FPD remains aligned with downstream performance as labeling schemes change and the sampling budget varies. Specifically, we evaluate FPD under alternative wind-power ramp labeling schemes and across multiple sample budgets.
We use the SDWPF dataset [24] with 10-minute resolution and reshape each example as a 1-hour window of six 10-minute values per turbine ($X\in\mathbb{R}^{N\times 6}$). Let the ramp rate be
$r=\frac{P_{50}-P_{00}}{P_{\max}},$
where $P_{00}$ and $P_{50}$ are the first and last 10-minute values within the hour.
We label SDWPF ramp events by direction and magnitude using two threshold sets inspired by [25] and [26] as shown in Table I.
TABLE I: Ramp-rate category definitions for the two labeling scenarios.
Category Scenario 1 (coarser) Scenario 2 (finer)
Strong down-ramp r<0.50r<-0.50 r<0.30r<-0.30
Moderate down-ramp $0.50\leq r<-0.33$ $0.30\leq r<-0.20$
Mild down-ramp $0.33\leq r<-0.25$ $0.20\leq r<-0.10$
Neutral $0.25\leq r\leq 0.25$ $0.10\leq r\leq 0.10$
Mild up-ramp $0.25<r\leq 0.33$ $0.10<r\leq 0.20$
Moderate up-ramp $0.33<r\leq 0.50$ $0.20<r\leq 0.30$
Strong up-ramp $r>0.50$ $r>0.30$
We then train two Wasserstein GAN (W-GAN) generators: W-GAN-full on the complete dataset (all categories) and W-GAN-collapse after removing all up-ramping categories to induce mode collapse (the model learns only down-ramp patterns). For each scenario, an identical MLP classifier is trained on real data labeled by that scenario and applied to synthetic samples from each generator at three sizes (0.2k/2k/20k). We report weighted-F1 and macro-F1. Weighted-F1 is the weighted average F1 score, so it reflects performance on prevalent categories and is a proxy for the perceived quality of typical generations. Macro-F1 is the unweighted mean of the F1 score, so it treats all categories equally and is sensitive to diversity. As shown in Table II, both generators achieve similarly high weighted F1, which indicates high-quality generation on prevalent categories, whereas macro-F1 drops markedly for W-GAN-collapse, revealing it is ignoring up-ramp categories, which is a classic mode collapse.
To compute the FPD score, we first upsample the real and synthetic series to 5-minute resolution, extract hourly features, and then measure the Fréchet distance between the generated and real feature distributions. Across both scenarios and all sample sizes, W-GAN-full consistently attains lower (better) FPD than W-GAN-collapse, with values stable as sample size increases. These results demonstrate that FPD aligns with the downstream classification diagnosis and remains robust to the choice of labeling scenario as well as sampling size.
TABLE II: Classification and FPD (scaled by $10^{-4}$) results under two scenarios
Scenario 1 Scenario 2
W-GAN-full W-GAN-collapse W-GAN-full W-GAN-collapse
Weighted F1 Macro F1 FPD Weighted F1 Macro F1 FPD Weighted F1 Macro F1 FPD Weighted F1 Macro F1 FPD
Sample_0.2k 0.994 0.955 454.4 0.987 0.640 903.5 0.995 0.936 454.4 0.989 0.679 794.3
Sample_2k 0.998 0.984 549.1 0.980 0.602 923.3 0.994 0.918 549.1 0.989 0.673 731.7
Sample_20k 0.997 0.980 550.7 0.984 0.631 885.8 0.994 0.946 550.7 0.990 0.676 800.5
IV-E Cross-model comparison with FPD